OpenStudy (anonymous):

write as a single fraction 1+(3)/x+2

OpenStudy (callisto):

Is the question \[1+\frac{3}{x+2}\]?

OpenStudy (anonymous):

yeah

OpenStudy (callisto):

\[1+\frac{3}{x+2}\]\[=\frac{x+2}{x+2}+\frac{3}{x+2}\]\[=\frac{(x+2)+3}{x+2}\]\[=...?\]

OpenStudy (anonymous):

(x+5)/x+2. but cant that simplify into 5/2 thusly making it 2 (1)/2

OpenStudy (anonymous):

?

OpenStudy (callisto):

Nope, you cannot :|

OpenStudy (callisto):

You can't cancel the unknowns like that..

OpenStudy (anonymous):

so itd just be (x+5)/x+2?

OpenStudy (callisto):

Yup!

OpenStudy (anonymous):

sweet!!

OpenStudy (anonymous):

@Callisto, but isn't this correct?: \(\large \frac{x+5}{x+5}=\frac{\cancel x+5}{\cancel x+5}=\frac{5}{5}=1 \) ???? :)

OpenStudy (callisto):

Of course not. Shouldn't it be: \[\large \frac{x+5}{x+5}=\frac{\cancel {x+5}}{\cancel {x+5}}=1\]???